# Pentagram maps over rings, Grassmannians, and skewers

### v1

@@ 4 Polygons and pentagram maps over rings
In this section, we define the notion of a polygon in the projective plane over a ring (see Definition 4.2), as well as show that the pentagram map on polygons is well-defined (see Section 4.2). Furthermore, we show that, for matrix rings and dual numbers, the general notion of a pentagram map over a ring specializes to Grassmann and skewer pentagram maps respectively (Examples 4.8 and 4.10).
@@ 4.1 Polygons in projective spaces over rings
Over a field, the pentagram map is well-defined on generic polygons. In particular, one can always apply the pentagram to a polygon such that all consecutive triple of vertices are in general position. The general position requirement means that the points are not collinear. In the context of (stably finite) rings, the right condition that ensures that the pentagram map is defined is the spanning condition.
@@ 4.2 Pentagram maps over stably finite rings, Grassmannians, and skewers
Applying Lemma 4.6 to vertices $P_{i},P_{i+2},P_{i+1},P_{i+3}$ (in that specific order) of a polygon $P_{i}$, we get that the pentagram map on polygons is well-defined. We note that the pentagram map is not a map from the set of polygons to itself, since the image of a polygon under the pentagram map is not necessarily a polygon. Instead, it should be understood as a binary relation on the set of polygons. In the case when the base ring is a field, the pentagram map is, actually, a rational mapping.
@@ 5.1 Difference operators with coefficients in a ring
The ring of (left) difference operators over $R$ is a subring $R^{\mathbb{Z}}[T]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by operators of left multiplication by elements of $R^{\mathbb{Z}}$, along with the left shift operator $T\in\mathrm{End}_{R}(R^{\mathbb{Z}})$, given by $(Ta)_{i}:=a_{i+1}$. In other words, a difference operator is an $R$-endomorphism $\mathcal{D}:R^{\mathbb{Z}}\to R^{\mathbb{Z}}$ given by
@@
Let $M$ be a free finite-rank left $R$-module with basis $e_{1},\dots,e_{m}$. For
@@
which is 00 if and only if $\mathcal{D}(a_{i})$ for all $i$, which is the same as to say $\mathrm{Span}\,v\subset\mathrm{Ker}\,\,\mathcal{D}$. ∎
@@ 5.2 The correspondence between difference operators and polygons
In order to prove Theorem 5.7, we will first build a map from the space of all degree $d$ properly bounded left difference operators with coefficients in a stably finite ring $R$ to the space of polygons, up to projective equivalence, in the left projective space $\mathbb{P}^{d-1}(R)$. Then, we will consider difference operators up to the action bi-infinite sequences of units on left and right and show that our original map factors through this quotient. Finally, we will prove that the map on the quotient is a bijection.
Let us denote the space of all degree $d$ properly bounded difference operators over $R$ as $R^{\mathbb{Z}}[T]_{d}^{pb}$ and let us denote the space of polygons in $\mathbb{P}^{d-1}(R)$ up to projective equivalence as $\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Our first aim to build a map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$.
@@
So, $\lambda_{c}^{*}$ maps $\varepsilon_{\mathcal{D}c,i}\in(\mathrm{Ker}\,\mathcal{D}c)^{*}$ to a scalar multiple of $\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,\mathcal{D})^{*}$. Therefore, the associated projective transformation $\mathbb{P}((\mathrm{Ker}\,\mathcal{D}c)^{*})\to\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})$ maps the $R\varepsilon_{\mathcal{D}c,i}$ to $R\varepsilon_{\mathcal{D},i}$, as needed.∎
@@
So, $\psi^{*}$ is left multiplication by $c$, and $\mathrm{Ker}\,\mathcal{D}^{\prime}=c\mathrm{Ker}\,\mathcal{D}$. Therefore, $\mathrm{Ker}\,\mathcal{D}^{\prime}c=\mathrm{Ker}\,\mathcal{D}$, and, by the above lemma, we have $\mathcal{D}^{\prime}c=c^{\prime}\mathcal{D}$, as needed. ∎
@@
Suppose $v_{i}\in M$ are such that the sequence $P_{i}=Rv_{i}$ is a polygon. Then there exists a properly bounded degree $d$ difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ such that $\mathcal{D}v=0$.
@@
$\mathcal{D}:=\alpha_{0}+\dots+\alpha_{d-1}T^{i+d-1}-T^{d},$
we have $\mathcal{D}v=0$. Further, for any $i\in\mathbb{Z}$ we have
@@
and thus $\alpha_{0,i}$ is left-invertible and hence a unit for any $i\in\mathbb{Z}$. Therefore, $\mathcal{D}$ is properly bounded, as desired. ∎
The following says that the difference operator $\mathcal{D}$ from Proposition 5.13 provides a right inverse for the map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Thus, this map is surjective, and the proof of Theorem 5.7 is complete.
@@
Since $\mathcal{D}v=0$, and the list $v_{i},\dots,v_{i+d-1}$ is a basis in $M$ for all $i$, by Proposition 5.6 we have $\mathrm{Span}\,v=\mathrm{Ker}\,\mathcal{D}$. This means that the image of the mapping $\psi\colon M^{*}\to\mathbb{R}^{\mathbb{Z}}$ given by $\xi\mapsto(\xi(v_{i}))$ is precisely $\mathrm{Ker}\,\mathcal{D}$. Furthermore, since $R$ is stably finite, and $M^{*}$, $\mathrm{Ker}\,\mathcal{D}$ are free modules of the same rank, it follows that $\psi\colon M^{*}\to\mathrm{Ker}\,\mathcal{D}$ is a right $R$-module isomorphism. Consider the dual $\psi^{*}\colon(\mathrm{Ker}\,\mathcal{D})^{*}\to M$. It is a left $R$-module isomorphism. Further, for any $\xi\in M^{*}$, we have
@@
so $\psi^{*}(\varepsilon_{\mathcal{D},i})=v_{i}$, meaning that the associated projective map $\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})\to\mathbb{P}(M)$ takes the polygon associated with $\mathcal{D}$ to the polygon $Rv_{i}$, as needed. ∎
@@ 6.1 Pentagram maps in terms of difference operators
By Theorem 5.7, polygons in the (left) projective plane $\mathbb{P}^{2}(R)$, considered up to projective transformation, are in one-to-one correspondence with elements of the set $R^{\mathbb{Z}}[T]_{3}^{pb}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})$ of properly bounded left difference operators of degree three, up to left-right action. Accordingly, one can describe the pentagram map as a dynamical system on difference operators. In this section we give an intrinsic description of that dynamical system. For technical reasons, we consider the inverse pentagram map, rather than the pentagram map itself. It turns out that that this inverse, phrased in terms of difference operators, is described by a simple linear equation. In the case when the ground ring is a field, this result was established in [8].
@@
We note that if $\mathcal{D}\sim\widehat{\mathcal{D}}$, then $\mathcal{D}\sim\alpha\widehat{\mathcal{D}}$ for any $\alpha\in(R^{\mathbb{Z}})^{*}$. Therefore, we have a left action of $(R^{\mathbb{Z}})^{*}$ on $R^{\mathbb{Z}}[T]_{3}^{pb}$.
@@
    The operator $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$ exists and is unique up to the left action of $(R^{\mathbb{Z}})^{*}$ on $R^{\mathbb{Z}}[T]_{3}^{pb}$.
@@
Suppose that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{reg}$ is regular. We will show that $\Psi(\bar{\varphi}(\mathcal{D}))=\varphi(\psi(\mathcal{D}))$. This will establish the inclusion (6.3), as well as commutativity of the diagram (6.4). Let $v_{i}:=\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,\mathcal{D})^{*}$. Then, by construction of the map $\bar{\varphi}\colon R^{\mathbb{Z}}[T]_{3}^{pb}\to\mathfrak{P}(\mathbb{P}^{2}(R))$, the polygon $(Rv_{i})$ represents the class of polygons $\bar{\varphi}(\mathcal{D})$ associated to $\mathcal{D}$. Furthermore, by Proposition 5.8, we have that $\mathrm{Span}\,v=\mathrm{Ker}\,\mathcal{D}$. Therefore,
@@ 6.2 Pseudo-difference operators
Any pseudo-difference operator $\mathcal{Q}\in R^{\mathbb{Z}}((T))$ can be uniquely represented as a sum of a series $\mathcal{Q}=\sum_{i=m}^{\infty}\alpha_{i}T^{i}$ where $\alpha_{i}\in R^{\mathbb{Z}}$, $m\in\mathbb{Z}$, and convergence is relative to the non-Archimedean norm.
@@
Let $\mathcal{Q}=\sum_{i=m}^{\infty}\alpha_{i}T^{i}$. If $\alpha_{m}$ is a unit in $R^{\mathbb{Z}}$, then $\mathcal{Q}$ a unit in $R^{\mathbb{Z}}((T))$.
@@ 6.3 Lax representation for pentagram maps over rings
Given a projective equivalence class of polygons in $\mathbb{P}^{2}(R)$, consider the associated properly bounded difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{pb}$, and let $\mathcal{Q}:=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}\in R^{\mathbb{Z}}((T))$ (this is well-defined by Corollary 6.8). Note that since $\mathcal{D}$ is defined up to the left-right action of $(R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}$, the operator $\mathcal{Q}$ is defined up to the conjugation action of $(R^{\mathbb{Z}})^{*}$ on $R^{\mathbb{Z}}((T))$. Hence, we have a map $\mathfrak{P}(\mathbb{P}^{2}(R))\to R^{\mathbb{Z}}((T))/(R^{\mathbb{Z}})^{*},$ taking a projective equivalence class of polygons to the conjugacy class of the associated operator $\mathcal{D}_{+}^{-1}\mathcal{D}_{-}$. We claim that this map provides a Lax representation of the pentagram map in the sense polygons related by the pentagram map are being mapped to conjugate elements of $R^{\mathbb{Z}}((T))/(R^{\mathbb{Z}})^{*}$.
@@
$\tilde{\mathcal{Q}}=\tilde{\mathcal{D}}_{+}^{-1}\tilde{\mathcal{D}}_{-}=\tilde{\mathcal{D}}_{+}^{-1}\mathcal{Q}\tilde{\mathcal{D}}_{+},$
@@ 7 Invariants of pentagram maps on closed polygons
Here we use the Lax form found in the previous section to give invariants (i.e., conserved quantities, or first integrals) of the pentagram map on closed polygons (i.e., polygons with periodic vertices). In the case of a field, these invariants are known to form a maximal Poisson-commuting family [14]. We expect that similar results can be obtained for much bigger classes of rings.

### v2

@@ 4 Polygons and pentagram maps over rings
In this section, we define the notion of a polygon in the projective plane over a ring (see Definition 4.2), as well as show that the pentagram map on polygons is well-defined (Section 4.2). Furthermore, we show that, for matrix rings and dual numbers, the general notion of a pentagram map over a ring specializes to Grassmann and skewer pentagram maps respectively (Examples 4.8 and 4.10).
@@ 4.1 Polygons in projective spaces over rings
Over a field, the pentagram map is well-defined on generic polygons. In particular, one can always apply the pentagram to a polygon such that all consecutive triples of vertices are in general position. The general position requirement means that the points are not collinear. In the context of (stably finite) rings, the right condition that ensures that the pentagram map is defined is the spanning condition.
@@ 4.2 Pentagram maps over stably finite rings, Grassmannians, and skewers
Applying Lemma 4.6 to vertices $P_{i},P_{i+2},P_{i+1},P_{i+3}$ (in that specific order) of a polygon $P_{i}$, we get that the pentagram map on polygons is well-defined. We note that the pentagram map is not a map from the set of polygons to itself, since the image of a polygon under the pentagram map is not necessarily a polygon. Instead, it should be understood as a binary relation on the set of polygons. In the case when the base ring is a field, the pentagram map is, actually, a rational mapping (for explicit formulas see, e.g., [15]).
@@ 5.1 Difference operators with coefficients in a ring
The ring of (left) difference operators over $R$ is a subring $R^{\mathbb{Z}}[T]\subset\mathrm{End}_{R}(R^{\mathbb{Z}})$ generated by operators of left multiplication by elements of $R^{\mathbb{Z}}$, along with the left shift operator $T\in\mathrm{End}_{R}(R^{\mathbb{Z}})$, given by $(Ta)_{i}:=a_{i+1}$. In other words, a difference operator is an endomorphism $\mathcal{D}:R^{\mathbb{Z}}\to R^{\mathbb{Z}}$ of right $R$-modules given by
@@
Let $e_{1},\dots,e_{m}\in M$ be a basis. For
@@
which is 00 if and only if $\mathcal{D}(a_{i}){\,=0}$ for all $i$, which is the same as to say $\mathrm{Span}\,v\subset\mathrm{Ker}\,\,\mathcal{D}$. ∎
@@ 5.2 The correspondence between difference operators and polygons
In order to prove Theorem 5.7, we will first build a map from the space of all degree $d$ properly bounded left difference operators with coefficients in a stably finite ring $R$ to the space of polygons, up to projective equivalence, in the left projective space $\mathbb{P}^{d-1}(R)$. Then, we will consider difference operators up to the action of bi-infinite sequences of units on left and right and show that our original map factors through this quotient. Finally, we will prove that the map on the quotient is a bijection.
Let us denote the space of all degree $d$ properly bounded difference operators over $R$ as $R^{\mathbb{Z}}[T]_{d}^{pb}$ and the space of polygons in $\mathbb{P}^{d-1}(R)$ up to projective equivalence as $\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Our first aim to build a map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$.
@@
So, $\lambda_{c}^{*}$ maps $\varepsilon_{\mathcal{D}c,i}\in(\mathrm{Ker}\,\mathcal{D}c)^{*}$ to a scalar multiple of $\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,\mathcal{D})^{*}$. Therefore, the associated projective transformation $\mathbb{P}((\mathrm{Ker}\,\mathcal{D}c)^{*})\to\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})$ maps $R\varepsilon_{\mathcal{D}c,i}$ to $R\varepsilon_{\mathcal{D},i}$, as needed.∎
@@
So, $\psi^{*}$ is left multiplication by $c$, and $\mathrm{Ker}\,\mathcal{D}^{\prime}=c\mathrm{Ker}\,\mathcal{D}$. Therefore, $\mathrm{Ker}\,\mathcal{D}^{\prime}c=\mathrm{Ker}\,\mathcal{D}$, and, by Lemma 5.11, we have $\mathcal{D}^{\prime}c=c^{\prime}\mathcal{D}$, as needed. ∎
@@
Suppose $v_{i}\in M$ are such that the sequence $P_{i}=Rv_{i}$ is a polygon. Then there exists a properly bounded degree $d$ difference operator $\mathcal{D}{{}_{v}}\in R^{\mathbb{Z}}[T]_{d}^{pb}$ such that $\mathcal{D}{{}_{v}}v=0$.
@@
$\mathcal{D}{{}_{v}}:=\alpha_{0}+\dots+\alpha_{d-1}T^{i+d-1}-T^{d},$
we have $\mathcal{D}{{}_{v}}v=0$. Further, for any $i\in\mathbb{Z}$ we have
@@
and thus $\alpha_{0,i}$ is left-invertible and hence a unit for any $i\in\mathbb{Z}$. Therefore, $\mathcal{D}{{}_{v}}$ is properly bounded, as desired. ∎
The following says that the correspondence $v\mapsto\mathcal{D}_{v}$ from Proposition 5.13 provides a right inverse for the map $\bar{\varphi}:R^{\mathbb{Z}}[T]_{d}^{pb}\to\mathfrak{P}(\mathbb{P}^{d-1}(R))$. Thus, this map is surjective, and the proof of Theorem 5.7 is complete.
@@
Since $\mathcal{D}v=0$, and the list $v_{i},\dots,v_{i+d-1}$ is a basis in $M$ for all $i$, by Proposition 5.6 we have $\mathrm{Span}\,v=\mathrm{Ker}\,\mathcal{D}$. This means that the image of the mapping $\psi\colon M^{*}\to\mathbb{R}^{\mathbb{Z}}$ given by $\xi\mapsto(\xi(v_{i}))$ is precisely $\mathrm{Ker}\,\mathcal{D}$. Furthermore, since $M^{*}$ and $\mathrm{Ker}\,\mathcal{D}$ are free modules of the same rank, and $R$ is stably finite, it follows that $\psi\colon M^{*}\to\mathrm{Ker}\,\mathcal{D}$ is a right $R$-module isomorphism. Consider the dual $\psi^{*}\colon(\mathrm{Ker}\,\mathcal{D})^{*}\to M$. It is a left $R$-module isomorphism, and, for any $\xi\in M^{*}$, we have
@@
so $\psi^{*}(\varepsilon_{\mathcal{D},i})=v_{i}$. Therefore, the associated projective map $\mathbb{P}((\mathrm{Ker}\,\mathcal{D})^{*})\to\mathbb{P}(M)$ takes the polygon associated with $\mathcal{D}$ to the polygon $Rv_{i}$, as needed. ∎
@@ 6.1 Pentagram maps in terms of difference operators
By Theorem 5.7, polygons in the (left) projective plane $\mathbb{P}^{2}(R)$, considered up to projective transformation, are in one-to-one correspondence with elements of the set $R^{\mathbb{Z}}[T]_{3}^{pb}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*})$ of properly bounded left difference operators of degree three, up to left-right action of $(R^{\mathbb{Z}})^{*}$. Accordingly, one can describe the pentagram map as a dynamical system on difference operators. In this section we give an intrinsic description of that dynamical system. For technical reasons, we consider the inverse pentagram map, rather than the pentagram map itself. It turns out that that this inverse, phrased in terms of difference operators, is described by a simple linear equation. In the case when the ground ring is a field, this result was established in [8].
@@
    The operator $\widehat{\mathcal{D}}\in R^{\mathbb{Z}}[T]_{3}^{pb}$ such that $\mathcal{D}\sim\widehat{\mathcal{D}}$ exists and is unique up to the left action of $(R^{\mathbb{Z}})^{*}$ on $R^{\mathbb{Z}}[T]_{3}^{pb}$. (Note that, if $\mathcal{D}\sim\widehat{\mathcal{D}}$, then $\mathcal{D}\sim\alpha\widehat{\mathcal{D}}$ for any $\alpha\in(R^{\mathbb{Z}})^{*}$. Therefore, the relation $\mathcal{D}\sim\widehat{\mathcal{D}}$ is invariant under the left action of $(R^{\mathbb{Z}})^{*}$ on $\widehat{\mathcal{D}}$.)
@@
Suppose that $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{reg}$ is regular. We will show that $\Psi(\bar{\varphi}(\mathcal{D}))=\varphi(\psi(\mathcal{D}))$. This will establish the inclusion $\varphi(R^{\mathbb{Z}}[T]_{3}^{reg}/((R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}))\subset\mathfrak{P}(\mathbb{P}^{2}(R))^{reg}$, as well as commutativity of the diagram (6.4). Let $v_{i}:=\varepsilon_{\mathcal{D},i}\in(\mathrm{Ker}\,\mathcal{D})^{*}$. Then, by construction of the map $\bar{\varphi}\colon R^{\mathbb{Z}}[T]_{3}^{pb}\to\mathfrak{P}(\mathbb{P}^{2}(R))$, the polygon $(Rv_{i})$ represents the class of polygons $\bar{\varphi}(\mathcal{D})$ associated to $\mathcal{D}$. Furthermore, by Proposition 5.8, we have that $\mathrm{Span}\,v=\mathrm{Ker}\,\mathcal{D}$. Therefore,
@@ 6.2 Pseudo-difference operators
Any pseudo-difference operator $\mathcal{Q}\in R^{\mathbb{Z}}((T))$ can be uniquely represented as a sum of a series $\mathcal{Q}=\sum_{i=m}^{\infty}\alpha_{i}T^{i}$ where $\alpha_{i}\in R^{\mathbb{Z}}$, $m\in\mathbb{Z}$, and convergence is relative to the above norm.
@@
Let $\mathcal{Q}=\sum_{i=m}^{\infty}\alpha_{i}T^{i}$. If $\alpha_{m}$ is a unit in $R^{\mathbb{Z}}$, then $\mathcal{Q}$ is a unit in $R^{\mathbb{Z}}((T))$.
@@ 6.3 Lax representation for pentagram maps over rings
Given a projective equivalence class of polygons in $\mathbb{P}^{2}(R)$, consider the associated properly bounded difference operator $\mathcal{D}\in R^{\mathbb{Z}}[T]_{3}^{pb}$, and let $\mathcal{Q}:=\mathcal{D}_{+}^{-1}\mathcal{D}_{-}\in R^{\mathbb{Z}}((T))$ (this is well-defined by Corollary 6.8). Note that since $\mathcal{D}$ is defined up to the left-right action of $(R^{\mathbb{Z}})^{*}\times(R^{\mathbb{Z}})^{*}$, the operator $\mathcal{Q}$ is defined up to the conjugation action of $(R^{\mathbb{Z}})^{*}$ on $R^{\mathbb{Z}}((T))$. Hence, we have a map $\mathfrak{P}(\mathbb{P}^{2}(R))\to R^{\mathbb{Z}}((T))/(R^{\mathbb{Z}})^{*},$ taking a projective equivalence class of polygons to the conjugacy class of the associated operator $\mathcal{D}_{+}^{-1}\mathcal{D}_{-}$. We claim that this map provides a Lax representation of the pentagram map in the sense that polygons related by the pentagram map are being mapped to conjugate elements of $R^{\mathbb{Z}}((T))/(R^{\mathbb{Z}})^{*}$.
@@
$\tilde{\mathcal{Q}}=\tilde{\mathcal{D}}_{+}^{-1}\tilde{\mathcal{D}}_{-}=\tilde{\mathcal{D}}_{+}^{-1}\mathcal{Q}\tilde{\mathcal{D}}_{+},$ (6.5)
@@
We note that, when $R$ is a field, (6.5) coincides with the Lax representation of the classical pentagram map given in [6, 8].
@@ 7 Invariants of pentagram maps on closed polygons
Here we use the Lax form found in the previous section to give invariants (i.e., conserved quantities, or first integrals) of the pentagram map on closed polygons (i.e., polygons given by periodic sequences $P_{i}\in\mathbb{P}^{2}(R)$). In the case of a field, these invariants are known to form a maximal Poisson-commuting family [15]. We expect that similar results can be obtained for much bigger classes of rings.